# A comparison of simulation tools for Muon-Induced X-ray Emission (MIXE) in thin films: a study case with lithium batteries

### v1

@@ 3.1 Negative muon implantation depths
The discrepancies in Tab. LABEL:tab:depths stem from SRIM’s averaging procedure. When a large low-density region precedes the target (e.g., air), the mean depth is dominated by those trajectories. Fitting the implantation tail with a Gaussian yields more realistic ranges consistent with PHITS, as shown in Fig. 4. The limited floating-point precision (5 digits, fixed µm unit) also restricts SRIM’s spatial resolution to $\sim$10 µm over a 10 cm setup.
@@
Figure 4: 20 MeV/c muon implantation profile as computed by SRIM (blue) for a 10 cm air cell (light blue) before the battery. The Gaussian fit (dashed line) yields 100218 µm ($R^{2}$ = 0.97), while SRIM’s internal mean is 97570 µm. PHITS (red) highlights the density transition at the first PE layer (light grey).
@@
Despite these limitations, SRIM remains useful for preliminary range estimates. For subsequent analyses, only GEANT4 and PHITS were retained. The mean implantation depths from both codes are shown in Fig. 5. Excellent agreement (<1%) is achieved irrespective of PHITS’ nspred flag, indicating limited sensitivity to angular straggling. Momentum values between 23–28 MeV/c were thus selected for optimal probing between the Al and Cu electrodes. The residuals between codes (Fig. 6) show deviations only at low momentum, dominated by scattering in upstream layers.
@@
Figure 5: Simulated mean implantation depths of negative muons in the battery cell. Dashed lines indicate layer interfaces. Vertical markers translate the standard deviations of the mean implantation depth.
@@
Figure 6: Solid lines: relative differences in mean muon implantation depth between PHITS and GEANT4 (reference). Larger deviations occur at low momentum due to scattering in air. Dashed lines: relative differences in standard deviation of the mean implantation depth as function of momentum, between PHITS and GEANT4 (reference).
@@
The implantation profiles for several momenta (Fig. 7) illustrate these effects. Anomalous discontinuities—e.g., around 425 µm in the Whatman layer at 27.4 MeV/c—appear only with nspred=1 and result from numerical inconsistencies in the slowing-down calculations of PHITS.
@@
Figure 7: Implantation profiles of negative muon as a function of beam momentum. Results are presented with the GEANT4 results as a reference (dashed line) and for PHITS having the physics card nspred = 0 (dots) and nspred = 1 (solid line). Graphs are labeled with corresponding muon beam momenta.
3.2 X-ray emission from negative muon cascades
@@
PHITS-computed photon spectra exhibit typical MIXE lines (Fig. 8), with low-energy $K$ transitions (C, Si, Al) well reproduced, while heavier elements (Ni, Mn, Co) show energy-shifted peaks. This issue is systematic across datasets.
@@
Figure 8: Raw MIXE emission spectrum from PHITS at 23.8 MeV/c. Low-energy peaks correspond to light elements K-lines MIXE signal and heavier elements (Mn, Ni, Co) L-lines. High-energy peaks corresponding to K-lines of the latter show shifted positions compared to reference data.
@@
Comparison between PHITS (Akylas’ code) and MUDIRAC reveals an increasing energy shift for high-$Z$ elements (Fig. 9), reaching several hundred keV for Ni and Cu. The underlying cause was traced to an energy-scaling bug in aama.f routine called by PHITS to compute the muonic X-ray cascade. The relative shift is shown in Fig. 9, with detailed values in Tab. A.2. To pinpoint which peak is associated to each isotope, we ran single element test run with 16 different targets. A dump file, called by modifying PHITS executable, would output the detailed cascade generated by aamaa.f, including transition levels, energies and intensities. It is noteworthy to mention that all hydrogen atoms had to be swapped for helium ones to avoid crashing of the aama.f routine, translating a pointer addressing issue in the atomic database. The slight modification of muon stopping power in virtual helium polymers has little impact on the scope of this assessment.
@@
Figure 9: Relative energy shift between PHITS and MUDIRAC for K and L transitions. K-lines are more affected than L-lines, with the discrepancy increasing with $Z$.
After identifying these systematic offsets, the relevant peaks were isolated in the composite spectrum (Fig. 10). The evolution of MIXE peaks with momentum reflects how the muons progressively penetrate deeper into the battery layers.
Figure 10: MIXE signal as estimated by PHITS for various momenta in the test battery. X-ray intensities are reported in Arbitrary Units, energies in keV. Each layer constitutive element should appear at a given momenta/implantation depth, such as estimated on Fig.7.
At low momentum, only Al lines are visible, followed by Ni, Mn, Co, and Si as muons penetrate deeper. Cu K-lines emerge only at 28 MeV/c, in line with experimental findings. The evolution of K-line intensities with momentum is plotted in Fig. 11.
Figure 11: PHITS calculated K-line intensities (arbitrary units) versus beam momentum. Low-intensity lines are magnified by a factor of ten (dashed). Cu peaks appear only at the highest momenta.
@@
While PHITS reproduces general spectral trends, absolute energies deviate from MUDIRAC, particularly for Ni and Cu isotopes (Tab. LABEL:tab:iso_shift). MUDIRAC solves the Dirac equation for a bound muon in a finite-size nucleus using realistic nuclear charge distributions (typically Fermi or homogeneously charged spheres), and includes vacuum-polarization and relativistic recoil corrections at leading order. The code provides transition energies with sub-keV accuracy for medium- and high-Z nuclei, and its predictions have been extensively benchmarked against precision muonic-atom X-ray measurements. PHITS underestimates isotopic energy shifts by nearly an order of magnitude compared to MUDIRAC and measurements.
@@
Table 5: Energies in keV of some elements’ muonic K–L transitions, as computed by MUDIRAC and PHITS.
@@
The transition energies of nickel K-lines measured experimentally (Fig. 12) agree closely with the predictions of MUDIRAC. A Gaussian decomposition performed on K and L lines groups (Tab. LABEL:tab:Ni_fit) shows consistent energy spacing and intensity ratios across isotopes. The areas of these Gaussian curves will be used to estimate the ratio of intensities between K and L emission lines.
@@
The ratio $\textbf{R}=K{}_{2}/K{}_{1}=0.5$ from PHITS was used to constrain Gaussian amplitudes, resulting in excellent fits to corrected K-line data. The L-lines around 308 and 313 keV exhibit similar proportionality. From the ensemble of Gaussian fits on the experimental data, the intensity ratio of nickel isotopes for $K/L$ was calculated. It yields 3.72, while PHITS returns 1.59. This simulated value is encouraging, as few models have demonstrated a relative intensity ratio of the same order of magnitude as the measurements.
@@ 4 Conclusion
A detailed benchmarking of GEANT4, PHITS, and SRIM has been conducted for the simulation of Muon-Induced X-ray Emission in a multilayer lithium-ion battery cell. GEANT4 and PHITS yield consistent predictions for implantation depths of negative muons, even in geometries with pronounced density contrasts, confirming their suitability for quantitative MIXE analysis. SRIM, while not originally intended for muon transport, remains a useful and rapid estimator of stopping ranges within individual layers, though its precision deteriorates when extended low-density materials such as air or other gases precede the target.
The modeling of muon capture and atomic cascades in PHITS reproduces relative line intensities with acceptable margins, but exhibits systematic energy shifts in K-line transitions of medium- to high-Z elements. This limitation currently prevents direct spectral comparison, yet the consistency of the predicted intensities underscores the reliability of the underlying cascade model. A practical remedy would be to couple PHITS to tabulated muonic-atom transition energies—either via lookup tables generated with MUDIRAC or by directly integrating the MUDIRAC routine into the PHITS atomic-cascade module—while retaining PHITS’ legacy cascade logic to compute relative intensities. Such a hybrid approach would substantially improve predictive accuracy.
@@
Table A.2: Relative difference in energy between PHITS and MUDIRAC for various isotopes. Values are expressed as percentage deviations from MUDIRAC reference energies. PHITS values for heavy elements such as gold or lead could not be computed.

### v2

@@ 3.1.4 Effect of muon capture in upstream materials
The discrepancies in Tab. 2 stem from SRIM’s averaging procedure. When a large low-density region precedes the target (e.g., air), the mean depth is dominated by those trajectories. Fitting the stopping depth tail with a Gaussian yields more realistic ranges consistent with PHITS, as shown in Fig. 4. The limited floating-point precision (5 digits, fixed µm unit) also restricts SRIM’s spatial resolution to $\sim$10 µm over a 10 cm setup.
@@
Figure 4: 20 MeV/c muon stopping depth profile as computed by SRIM (blue) for a 10 cm air cell (light blue) before the battery. The Gaussian fit (dashed line) yields 100218 µm ($R^{2}$ = 0.97), while SRIM’s internal mean is 97570 µm. PHITS (red) highlights the density transition at the first PE layer (light grey).
Despite these limitations, SRIM remains useful for preliminary range estimates. Even without muon capture physics implemented, SRIM is capable of reproducing ranges of negative muons with an accuracy similar to PHITS with its extended muon physics routines, as seen in Tab.4. For subsequent analyses, only GEANT4 and PHITS were retained. The mean stopping depths and distribution spread from both codes are shown in Fig. 5.
3.1.5 Sensitivity to transport settings
@@
Agreement at the percent level is obtained irrespective of the PHITS nspred flag, indicating limited sensitivity to the corresponding angular straggling settings for this geometry. Momentum values between 23–28 MeV/c were thus selected for optimal probing between the Al and Cu electrodes. The residuals between codes (Fig. 6) show discrepancies only at low momentum, dominated by scattering in upstream layers.
@@
Figure 5: Simulated mean stopping depths of muons in the battery cell. Dashed lines indicate layer interfaces. Vertical markers translate the standard deviations of the mean stopping depth.
@@
Figure 6: Solid lines: relative differences in mean muon stopping depth between PHITS and GEANT4 (reference). Larger discrepancies occur at low momentum due to scattering in air. Dashed lines: relative differences in standard deviation of the mean stopping depth as function of momentum, between PHITS and GEANT4 (used as reference model).
@@
The stopping depth profiles for several momenta (Fig. 7) illustrate these effects. Anomalous discontinuities (e.g., around 425 µm in the Whatman layer at 27.4 MeV/c) appear only with nspred=1 and result from numerical inconsistencies in the slowing-down calculations of PHITS.
@@
Figure 7: stopping depth profiles of muon as a function of beam momentum. Results are presented with the GEANT4 results as a reference (dashed line) and for PHITS having the physics card nspred = 0 (dots) and nspred = 1 (solid line). Graphs are labeled with corresponding muon beam momenta. Muons come from the left, as represented in Fig. 3
3.2 X-ray emission from muon cascades
3.2.1 Spectral overview and transition energy discrepancies
@@
PHITS-computed photon spectra exhibit typical MIXE lines (Fig. 8), with low-energy $K$ transitions (C, Si, Al) well reproduced, while heavier elements (Ni, Mn, Co) show energy-biased peaks. This issue is systematic across datasets.
@@
Figure 8: Raw MIXE emission spectrum from PHITS at 23.8 MeV/c. Low-energy peaks correspond to light elements K-lines MIXE signal and heavier elements (Mn, Ni, Co) L-lines. High-energy peaks corresponding to K-lines of the latter show shifted positions compared to reference data. The error bars represent the one-standard-deviation (1 ) statistical uncertainties provided by PHITS for each energy bin. These correspond to the relative Monte Carlo tally error (r.err) multiplied by the bin flux value (1 keV wide), reflecting counting statistics from the finite number of simulated particle histories.
@@
Comparison between PHITS (Akylas’ code) and MuDirac reveals an increasing energy discrepancy for high-$Z$ elements (Fig. 9), reaching several hundred keV for Ni and Cu. In the present PHITS version, the dominant contribution to this bias was traced to an energy scaling issue in the aama.f cascade routine; more generally, low-$n$ (K-series) muonic levels are increasingly sensitive to relativistic and finite-nuclear-size corrections with increasing $Z$, which explains why such discrepancies tend to grow for heavier elements and affect K lines more strongly than L lines. The relative difference is shown in Fig. 9, with detailed values in Tab. A.2. To pinpoint which peak is associated to each isotope, we ran single element test run with 16 different targets. A dump file, called by modifying PHITS executable, would output the detailed cascade generated by aamaa.f, including transition levels, energies and intensities. To enable diagnostic dumping of the aama.f cascade routine, hydrogen-bearing materials were temporarily replaced by helium in a dedicated technical test configuration to avoid a known instability of the PHITS atomic database interface. This substitution was used only for debugging the cascade bookkeeping and does not affect the transport/comparison results presented above.
@@
Figure 9: Relative energy difference between PHITS and MuDirac for K and L transitions. K-lines are more affected than L-lines, with the discrepancy increasing with $Z$.
After identifying these systematic offsets, the relevant peaks were isolated in the composite spectrum, shown in Appendix. The evolution of MIXE peaks with momentum reflects how the muons progressively penetrate deeper into the battery layers. At low momentum, only Al lines are visible, followed by Ni, Mn, Co, and Si as muons penetrate deeper. Cu K-lines emerge only at 28 MeV/c because the muon stopping distribution reaches the Cu current collector located at the back of the multilayer stack; at lower momenta the stopping probability at that depth is negligible (see muon stopping depth profiles in Fig. 7). The evolution of K-line intensities with momentum is plotted in Fig. 10. Error bars represent the 1  statistical uncertainties of the PHITS Monte Carlo tallies. For each K-line, the uncertainty was obtained by propagating the per-bin relative error over a ±1 keV energy window. When multiple K-lines of the same element contribute at a given momentum, the plotted value corresponds to their mean intensity and the uncertainty is propagated assuming independent statistics. The quoted errors therefore reflect only counting (Monte Carlo) statistics.
[…]
@@
While PHITS reproduces general spectral trends, absolute energies deviate from MuDirac, particularly for Ni and Cu isotopes (Tab. LABEL:tbl:iso_shift). MuDirac solves the Dirac equation for a bound muon in a finite-size nucleus using realistic nuclear charge distributions (typically Fermi or homogeneously charged spheres), and includes vacuum-polarization and relativistic recoil corrections at leading order. The code provides transition energies with sub-keV accuracy for medium- and high-Z nuclei, and its predictions have been extensively benchmarked against precision muonic-atom X-ray measurements. PHITS underestimates isotopic energy shifts by nearly an order of magnitude compared to MuDirac and measurements.
@@
Table 5: Energies in keV of some elements’ muonic $\mathrm{K{}_{1}}$ and $\mathrm{K{}_{2}}$ transitions, as computed by MuDirac and PHITS.
@@
The transition energies of nickel K-lines measured experimentally (Fig. 11) agree closely with the predictions of MuDirac. A Gaussian decomposition performed on K and L lines groups (Tab. LABEL:tbl:Ni_fit) shows consistent energy spacing across isotopes. The areas of these Gaussian curves will be used to estimate the ratio of intensities between K and L emission lines.
Nickel was selected as a test case because its relatively high atomic number leads to intense and well-separated muonic K transitions at energies above 1.4 MeV. At these energies, self-absorption within the sample is negligible and attenuation in the surrounding materials is minimal compared to the L lines. In addition, the natural isotopic composition of Nickel produces clearly resolved isotope-dependent shifts, providing a stringent test of the cascade and transport modeling.
@@
3.2.2 Nickel spectral fit and relative intensities
The relative line intensities were extracted by modeling each transition with a Gaussian response function of fixed instrumental width , assuming detector resolution dominates over the intrinsic line width. For the K series, the amplitudes of the doublet components were constrained such that
$R_{K}=\frac{I(K{}_{2})}{I(K{}_{1})}=0.5,$ (3)
as obtained from PHITS and the Akylas cascade code. This constraint was imposed consistently for all Nickel isotopes, thereby reducing the number of free parameters and enforcing physical consistency of the fit. No analogous constraint was applied to the $\mathrm{L{}_{1}}$ and $\mathrm{L{}_{1}}$ lines around 308 and 313 keV, as no universal fixed ratio was observed across the set of assessed elements.
For each Nickel isotope $i$, the fitted peak area $A_{i,\ell}$ satisfies
[…]
The global intensity ratio between K and L emission was therefore defined as
$\frac{K}{L}=\frac{\sum_{i}\left[A_{i,K{}_{1}}+A_{i,K{}_{2}}\right]}{A_{L{}_{1}}+A_{L{}_{1}}}.$ (5)
It should be emphasized that the ratio defined above is not a purely atomic quantity. Because the detector efficiency, photon attenuation in the sample and surrounding materials, and the experimental geometry enter differently at 300 keV and 1.4 MeV, the observable $\mathrm{K/L}$ is inherently geometry dependent. The comparison therefore probes not only the atomic cascade yields $\mathrm{Y_{i,\ell}}$ but also the consistency of photon transport and efficiency treatment in the simulation. From the ensemble of Gaussian fits to the experimental spectrum retrieved from datasets of [edouard], compiled in Tab. LABEL:tbl:Ni_fit, this ratio is found to be
$\left(\frac{K}{L}\right)_{\mathrm{exp}}=3.72$ (6)
The corresponding value obtained from PHITS is
$\left(\frac{K}{L}\right)_{\mathrm{PHITS}}=1.54\pm 0.08.$ (7)
The quoted uncertainty reflects statistical propagation of the fitted peak-area uncertainties. The experimental ratio, however, is affected by additional systematic contributions that are not explicitly available in the published dataset. These include HPGe energy calibration and efficiency calibration uncertainties, geometrical alignment and solid-angle effects, dead-time corrections, as well as possible self-absorption and background-subtraction systematics. Consequently, the comparison should be interpreted primarily at the level of global consistency rather than as a precision benchmark.
Although PHITS underestimates the relative K-to-L emission strength, it reproduces the correct order of magnitude and preserves the expected hierarchy between high-energy and low-energy cascade transitions. Given the strong sensitivity of L-line intensities to atomic cascade modeling, low-energy attenuation, and detector efficiency corrections, this agreement provides a meaningful validation of the coupled cascade–transport treatment implemented in the code. Further refinement would require a fully efficiency-corrected experimental spectrum and a detailed treatment of low-energy photon transport in the experimental geometry.
@@ 4 Conclusion
A detailed benchmarking of GEANT4, PHITS, and SRIM has been conducted for the simulation of Muon-Induced X-ray Emission in a multilayer lithium-ion battery cell. GEANT4 and PHITS yield consistent predictions for stopping depths of muons, even in geometries with pronounced density contrasts, confirming their suitability for quantitative MIXE analysis. SRIM, while not originally intended for muon transport, remains a useful and rapid estimator of stopping ranges within individual layers, though its precision deteriorates when extended low-density materials such as air or other gases precede the target.
The modeling of muon capture and atomic cascades in PHITS reproduces relative line intensities with acceptable margins, but exhibits systematic energy discrepancies in K-line transitions of medium- to high-Z elements. This limitation currently prevents direct spectral comparison, yet the consistency of the predicted intensities underscores the reliability of the underlying cascade model. A practical remedy would be to couple PHITS to tabulated muonic-atom transition energies—either via lookup tables generated with MuDirac or by directly integrating the MuDirac routine into the PHITS atomic-cascade module—while retaining PHITS’ legacy cascade logic to compute relative intensities. Such a hybrid approach would substantially improve predictive accuracy.
@@
Table A.2: Relative difference in energy between PHITS and MuDirac for various isotopes. Values are expressed as percentage difference from MuDirac reference energies. PHITS values for heavy elements such as gold or lead could not be computed.
@@
Figure 12: MIXE signal as estimated by PHITS for various momenta in the test battery. Each layer constitutive element should appear at a given momenta/stopping depth, such as estimated on Fig. 7. Error bars are 1 Monte-Carlo statistical uncertainties from PHITS. A nominal 0.5 keV resolution ($\approx$ half a tally bin) is assumed to merge unresolved transitions and to define the energy integration window across adjacent bins.